If 50 cm3 of 0.1 moldm-3 HCl and 50 cm3 of 0.1 moldm-3 NaOH are mixed, the temperature of the solution rises by 0.68 oC. Calculate the enthalpy change of the reaction in kJmol-1.
1
Expert's answer
2020-08-23T07:44:23-0400

HCl and NaOH are Strong Acid and Strong Base So, Enthaply of neutralisation is -13.7 KCal per mol.


Reacting mol of HCl and NaOH are

50 × 0.1 / 1000 mol .


So, Enthaply Change =

- (13.7 × 5 × 10^-3 ) KCal. = 0.0685 KCal.


Enthalpy Change =( 0.0685 × 1000 )= -68.5 Cal.

Enthalpy change = - 68.5 Cal = 68.5 × 4.2 j.= - 287.7 j ANSWER